OpenStudy (anonymous):

Evaluate the integral 15^x dx on -1,1

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}15^{x}dx\]

OpenStudy (anonymous):

15^x=e^(log(15^x))=e^(log(15)x)

OpenStudy (australopithecus):

So, \[\int\limits{a^u}du = \frac{a^u}{\ln(a)} + c\]

OpenStudy (anonymous):

I've gotten as far as \[\frac{ 15^x }{\ln15 }\] But when I plug in -1 and 1 and subtract, my answer doesn't match.

OpenStudy (australopithecus):

Blah I forgot the notation but you need to subtract the lower limit from the upper limit

OpenStudy (australopithecus):

15/ln(15) - (15^-1)/ln(15)

OpenStudy (australopithecus):

Do you follow?

OpenStudy (anonymous):

Yes. That's what I was doing but the answer is supposed to be \[\frac{ 244 }{15\ln15}\] I can see how the bottom turns to 15ln15 but the 244 has be completely lost.

OpenStudy (australopithecus):

remember this is fractional subtraction

OpenStudy (australopithecus):

matching denominators and all that jazz

OpenStudy (australopithecus):

that is the same answer as what I posted above as well

OpenStudy (anonymous):

Ah, of course! It all makes sense now! Thank you so much for pointing that out!!

OpenStudy (australopithecus):

No problem anytime :D I hope your life is going well

OpenStudy (australopithecus):

and I just got your name on here ugh I'm ditzy

OpenStudy (australopithecus):

https://www.wolframalpha.com/input/?i=integrate+between+-1+and+1+for+15^x+dx If you didnt know how to integrate using wolfram alpha

OpenStudy (australopithecus):

Sometimes the answer will look different from what you got but in reality it is the same so I really recommend using wolfram alpha to save yourself the time

OpenStudy (anonymous):

I was trying to use it but couldn't figure out how to enter the bounds. Now I know!